# Separate Motion from Appearance: Customizing Motion via Customizing Text-to-Video Diffusion Models

### v1

@@ 3.3 Appearance Highway
The underlying logic of AH is that as the skip connection of U-Net mainly conveys high-frequency appearance information [28], the starting point alternation can provide a “highway” for the hidden states of non-adapted spatial transformers and cut the shortcut from the hidden states of LoRA-adapted temporal transformers. As shown in Fig. 3, the videos are generated conditioned on “A monkey is playing golf on a field full of flowers”. The first row of Fig. 3 illustrates the videos generated with AH. We further show how the appearance changes as we enhance the effect of hidden states from the spatial transformers by varying scales. Compared with adapted T2V-DM with vanilla skip connection, AH yields videos with appearance more aligned with text descriptions of the target. Moreover, as the effect of hidden states from the spatial transformers is enhanced by increasing scales (see first row of Fig. 3), the appearance of generated videos becomes more aligned with the text descriptions. In contrast, with vanilla skip connection, no matter how we enhance the temporal transformer’s output, the appearance cannot match the text descriptions.
@@
Figure 3: The comparison between the Appearance Highway and the Skip Connection is illustrated in the figure. The figure presents videos generated by “A monkey is playing golf on a field full of flowers”. The first row shows that as the scale of the appearance highway increases, appearance leakage is progressively alleviated. In contrast, the second row shows that increasing the scale of the vanilla skip connection has no impact on appearance leakage. The smaller images below each main image represent subsequent frames of the video.
@@
Figure 4: Investigation of appearance highway. (a) shows the similarity between the TAP/AH’s hidden states and T2V-DM’s hidden states. (b) shows the similarity between the hidden states of TAP/AH and hidden states of T2V-DM with inserted spatial LoRAs. (c) shows the accuracy and confidence of the motion classifier.
@@
Further, we find that AH may largely maintain the appearance generation ability of vanilla T2V-DM, without harming the motion concept learning from the reference videos. We conduct experiments to verify this. As changing the starting point of skip connection directly affects the decoding results, we extract the hidden states from the decoding part after projecting the hidden states combined from the starting point of skip connection and the decoder input (the pink part illustrated in Fig. 2(c)).
@@
We compare hidden states extracted from four kinds of models: (1) vanilla T2V-DM with high appearance generation ability; (2) T2V-DM with spatial-LoRA (SL) which customizes the appearance of reference videos; (3) T2V-DM with TAP (TAP); (4) TAP combined with AH (AH).
@@
As shown in Fig. 4 (a), we compare the similarity score between AH and T2V-DM and that between TAP and T2V-DM. We observe that with AH, the hidden states of decoder is more similar to T2V-DM compared to TAP, which indicates AH may further improves the appearance generation ability of adapted T2V-DM. We also compare the similarity score between AH and SL and that between TAP and SL in Fig. 4 (b). We observe that with AH, the hidden states of decoder is more dissimilar to SL compared to TAP. All those results indicate AH may enhance the appearance modeling ability beyond TAP.
@@
We extract optical flows from TAP’s hidden states as positive samples and those from T2V-DM’s hidden states as negative samples to train a motion classifier. Then we use this classifier to evaluate the hidden states from AH. Fig. 4 shows most of the hidden states from AH can be classified into TAP rather than vanilla T2V-DM in a high accuracy and with high confidence. These results show that utilizing AH may not disturb the motion generation ability of TAP.
@@
3.4 Phased LoRA Integration
Figure 5: Qualitative results on one-shot. Tune-A-Video [38], VMC [18], MotionDirector [47] and our method use single text-video pair from the LOVEU-TGVE-2023 dataset [39] as the reference video. VideoComposer [35] and Ctrl-A-Video [7] only need single video. Then we use an editing prompt with a different appearance description to generate a new video.
@@ 3.4 Phased LoRA Integration
Methods Text Align. Temporal Consist. Aesthetic Score ViCLIP Score VBench Average Appearance Diversity Motion Fidelity Temporal Consistency
VideoComposer [35] 27.34 90.00 20.29 25.63 74.12 65.5/34.5 63.6/36.4 67.3/32.7
Control-A-Video [7] 25.56 93.40 19.80 23.66 73.33 74.5/25.5 58.2/41.8 58.2/41.8
Tune-A-Video [38] 25.69 92.11 20.18 23.37 75.34 58.2/41.8 56.4/43.6 60.0/40.0
VMC [18] 25.67 90.50 20.46 23.98 74.23 81.8/19.2 69.1/30.9 83.6/16.4
MotionDirector [47] 27.55 93.61 20.36 25.54 75.95 61.8/38.2 58.2/41.8 63.6/36.4
@@
Table 2: Quantitative results on one-shot. In human evaluation, the paired numbers shows our method’s voting rate on the left, while previous methods’ voting rate on the right.
Figure 6: Qualitative results on few-shot. During training, few videos from UCF Sports dataset [25, 30] are used as reference videos, with “person” as the main subject in prompts. During inference, we use Stable Diffusion to generate images paired with the texts, which serve as the image conditions for LAMP [40] and DreamVideo [37]. MotionDirector [47] and our method do not need image conditions.
@@
Methods Text Align. Temporal Consist. Aesthetic Score ViCLIP Score VBench Average Appearance Diversity Motion Fidelity Temporal Consistency
LAMP [40] 26.19 87.89 18.41 22.68 67.41 59.1/40.9 63.6/36.4 60.6/39.4
DreamVideo [37] 27.48 90.52 19.17 25.10 71.75 66.7/33.3 53.0/47.0 50.0/50.0
MotionDirector [47] 29.74 94.27 19.93 27.83 72.91 59.1/40.9 51.5/48.5 48.5/51.5
@@
Table 3: Quantitative results on few-shot. In human evaluation, the paired numbers shows our method’s voting rate on the left, while previous methods’ voting rate on the right.
@@
${z}_{t-1}=\begin{cases}\frac{1}{\sqrt{\alpha_{t}}}\left({z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\epsilon_{\theta}({z}_{t},t)\right)+\sigma_{t},&t\leq\tau\\ \frac{1}{\sqrt{\alpha_{t}}}\left({z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\epsilon_{\theta,\theta_{K}}({z}_{t},t)\right)+\sigma_{t},&t>\tau\end{cases}$ (3)
which $\alpha$ is a constant controlling the noise strength.
4 Experiment
4.1 Experimental Setup
@@
Datasets For evaluations on a single video, we first utilize LOVEU-TGVE-2023 dataset [39]. The dataset includes 76 reference videos, each accompanied by 4 editing prompts. Following  [47], we design 3 additional prompts with significant appearance changes for each reference video to evaluate the diversity of the generated videos. For evaluations on few videos, we select 3-5 reference videos from each of the 13 categories in the UCF Sports dataset [25, 30]. And we design 8 prompts for the corresponding reference videos for each category.
@@
Baselines For evaluations on one-shot, we compare our method with the popular motion customization method, including MotionDirector [47] and VMC [18]. In addition, we compare with controllable video generation methods, including VideoComposer [35], Control-A-Video [7], and Tune-a-Video [38]. Tune-a-Video, VMC, MotionDirector, and our method need fine-tuning, whereas VideoComposer and Control-A-Video do not need training. For evaluations on few-shot, we compare our approach with DreamVideo [37], LAMP [40], and MotionDirector.
@@
Metrics For automatic evaluations, we use 3 metrics: text alignment (Text Align.), temporal consistency (Temporal Consist.), and aesthetic score. Text alignment measures the average CLIP similarity [23] between each frame and the text prompt, reflecting how well the generated video aligns with the given description. Temporal consistency assesses the video’s coherence by CLIP similarity between frames. The aesthetic score is computed using an aesthetics model [20]. To provide a more comprehensive evaluation, we also leverage ViCLIP [36] and VBench [17]. ViCLIP is used as a supplementary metric alongside TextCLIP. VBench is a benchmark that evaluates models across multiple dimensions. Due to space limitations, we report the average of these results (VBench Average), with detailed results provided in Appendix. For human evaluations, we invite 11 participants to complete a survey, resulting in a total of 1,419 evaluation responses. The survey assesses appearance diversity, motion fidelity, and temporal consistency. We show two videos: one generated by our method and the other by previous methods. For each evaluation dimension, users vote for the better video.
4.2 Evaluations on One-Shot
@@
Quantitative results. For automatic evaluations in Tab. 2, our method significantly improves text alignment and ViCLIP score, demonstrating that it suppresses appearance leakage. Our method also demonstrates superiority in other dimensions. Human evaluations confirm the same results.
@@
Qualitative results. In the example on the left of Fig. 5, previous methods overfitted the “grey sharks”, while ours generate a realistic “sea dragons”. In the example on the right, our method generated a vividly “sandwich”, while the “sandwich” produced by other methods resembled the “pizza” from the reference video.
4.3 Evaluation on Few-Shot
@@
Quantitative results. As shown in 3, the automatic evaluation shows our method surpasses previous approaches across multiple dimensions. The human evaluations show that while our temporal consistency may not surpass that of DreamVideo and MotionDirector, our appearance diversity and motion fidelity are superior.
@@
Qualitative Results. Fig. 6 illustrates that MotionDirector occasionally exhibits appearance leakage from the training set. In the example on the left, DreamVideo generates a teddy bear holding an object resembling a suitcase and wearing a tie. It occurs because the training set includes humans carrying suitcases and wearing ties. MotionDirector generates a background featuring vertical reflective structures, similar to the background in the second reference video. In the right-hand case, the backgrounds generated by previous methods lack a distinctly “cyberpunk” aesthetic and instead resemble those of the reference videos. In contrast, our method effectively avoids this issue.
4.4 Ablation
@@
Effect of Our Sub-Contributions. in Tab. 4, we evaluate the capabilities of each contribution: temporal attention purification (TAP), appearance highway (AH), and phased LoRA integration (PLI). TAP, AH, and PLI each significantly enhance text alignment. AH improves text alignment and aesthetic score but slightly compromises temporal consistency. However, combined with PLI, it not only further enhances text alignment and aesthetic score but also mitigates the loss in motion consistency caused by AH.
Effect of motion customization with different T2V-DM. Tab. 4 demonstrates that our method is effective across various foundation models using a spatial-temporal U-Net. We select ZeroScope [31] and ModelScope [34] as the foundation models for verifying. The effectiveness of each component is confirmed in both models.
Training with AH vs. Post-processing with AH. The appearance highway can be used during both training and inference, or exclusively during inference. We compare the differences between these two usage methods. As shown in Tab. 5, the results of the two methods are comparable. For simplicity, we use it with post-processing as our method.
Models TAP AH PLI Text Align. Temporal Consist. Aesthetic Score ViCLIP Score VBench Average
@@
Table 4: Effect of each component of our method and effect of motion customization with different T2V-DM. We test on two popular T2V-DM models: ZeroScope(ZS) and ModelScope(MS).
Method Text Align. Temporal Consist. Aesthetic Score ViCLIP Score VBench Average
@@
Table 5: Training with AH vs. Post-possing with AH: (1) Training: apply it during both training and inference. (2) Post-processing: apply it only during inference. Experiments show that these two strategies are comparable.
5 Conclusion
In this paper, we propose two novel strategies to separate motion from appearance during adapting T2V-DM to perform motion customization task, including Temporal Attention Purification (TAP) and Appearance Highway (AH). In TAP, we utilize LoRA only to adapt the temporal attention and in AH, we alter the starting point of skip connection from LoRA-adapted temporal transformer block to non-adapted spatial transformer block. Extensive experiments show that our method performs favorably against previous state-of-the-arts.

### v2

@@ 3.3. Appearance Highway
As TAP may not perfectly avoid encoding appearance information in motion LoRAs, we propose using Appearance Highway (AH) strategy to maintain the appearance generation ability of the original T2V-DM and further avoid appearance leakage issues. Specifically, we alter the starting point of each skip connection in spatial-temporal U-Net (Çiçek et al., 2016; Ho et al., 2022b) from the output of each temporal transformer to the output of each spatial transformer. In other words, the appearance highway is an improved skip connection designed to decouple motion and appearance. The starting point alternation can provide a “highway” for the hidden-states of non-adapted spatial transformers and cut the shortcut from the hidden-states of LoRA-adapted temporal transformers. As shown in Fig. 3, the videos are generated conditioned on “A monkey is playing golf on a field full of flowers”. The first row of Fig. 3 illustrates the videos generated with AH. We further show how the appearance changes as we enhance the effect of hidden states from the spatial transformers by varying scales. Compared with adapted T2V-DM with vanilla skip connection, AH yields videos with appearance more aligned with text of the target. Moreover, as the effect of hidden states from the spatial transformers is enhanced by increasing scales (see the first row of Fig. 3), the appearance of generated videos becomes more aligned with the text descriptions. In contrast, with vanilla skip connection, no matter how we enhance the temporal transformer’s output, the appearance cannot match the text.
@@
Figure 3. The comparison between the Appearance Highway and the Skip Connection is illustrated in the figure. The figure presents videos generated by “A monkey is playing golf on a field full of flowers”. The first row shows that as the scale of the appearance highway increases, appearance leakage is progressively alleviated. In contrast, the second row shows that increasing the scale of the vanilla skip connection has no impact on appearance leakage. The smaller images below each main image represent subsequent frames of the video.
[…]
The comparison between the Appearance Highway and the Skip Connection is illustrated in the figure.
@@
Figure 4. Investigation of AH. (a) shows the similarity between the TAP/AH’s hidden states and T2V-DM’s. (b) shows the similarity between the hidden states of TAP/AH and hidden states of T2V-DM with inserted spatial LoRAs. (c) shows the accuracy and confidence of the motion classifier.
[…]
Investigation of AH.
@@
Furthermore, we find that AH may largely maintain the appearance generation ability of vanilla T2V-DM, without harming the motion concept learning from the reference videos. We conducted experiments to verify this. As changing the starting point of skip connection directly affects the decoding results, we extract the hidden states from the decoding part after projecting the hidden states combined from the starting point of skip connection and the decoder input (the pink part illustrated in Fig. 2(c)).
@@
We compare hidden states from 4 kinds of models: (1) T2V-DM with high appearance generation ability; (2) T2V-DM with spatial-LoRA (SL) which customizes the appearance of reference videos; (3) T2V-DM with TAP (TAP); (4) TAP combined with AH (AH).
@@
As shown in Fig. 4 (a), we compare the similarity between AH and T2V-DM and that between TAP and T2V-DM. We observe that with AH, the hidden states of decoder are more similar to T2V-DM, which indicates AH may further improve the appearance generation ability of adapted T2V-DM. We also compare the similarity between AH and SL and that between TAP and SL in Fig. 4 (b). We observe that with AH, the hidden states of the decoder are more dissimilar to SL compared to TAP. All of these results indicate that AH may enhance the modeling ability of the appearance beyond TAP.
@@
We extract optical flows from TAP’s hidden states as positive samples and those from T2V-DM’s hidden states as negative samples to train a motion classifier. Then we use this classifier to evaluate the hidden states from AH. Fig. 4 shows most hidden states from AH can be classified into TAP rather than vanilla T2V-DM in a high accuracy and with high confidence. These results show that utilizing AH may not disturb the motion generation ability of TAP.
@@
3.4. Phased LoRA Integration
Figure 5. Qualitative results on one-shot. Tune-A-Video (Wu et al., 2023a), VMC (Jeong et al., 2024), MotionDirector (Zhao et al., 2023) and our method need to fine-tune. VideoComposer (Wang et al., 2024) and Ctrl-A-Video (Chen et al., 2023) are training free.
[…]
Qualitative results on one-shot.
Table 2. Quantitative results on few-shot. In human evaluation, the paired numbers show our method’s voting rate on the left, while previous methods’ voting rate on the right.
@@ 3.4. Phased LoRA Integration
Methods Text Align. $\uparrow$ Temporal Consist. $\uparrow$ Aesthetic Score $\uparrow$ ViCLIP Score $\uparrow$ VBench Average $\uparrow$ Appearance Diversity Motion Fidelity Temporal Consistency
VideoComposer (Wang et al., 2024) 27.34 90.00 20.29 25.63 74.12 65.5/34.5 63.6/36.4 67.3/32.7
Control-A-Video (Chen et al., 2023) 25.56 93.40 19.80 23.66 73.33 74.5/25.5 58.2/41.8 58.2/41.8
Tune-A-Video (Wu et al., 2023a) 25.69 92.11 20.18 23.37 75.34 58.2/41.8 56.4/43.6 60.0/40.0
VMC (Jeong et al., 2024) 25.67 90.50 20.46 23.98 74.23 81.8/19.2 69.1/30.9 83.6/16.4
MotionDirector (Zhao et al., 2023) 27.55 93.61 20.36 25.54 75.95 61.8/38.2 58.2/41.8 63.6/36.4
@@
Figure 6. Qualitative results on few-shot. During training, few videos from UCF Sports dataset (Rodriguez et al., 2008; Soomro and Zamir, 2015) are used as reference videos. During inference, we use Stable Diffusion to generate images paired with the texts, which serve as the image conditions for LAMP (Wu et al., 2024) and DreamVideo (Wei et al., 2024). MotionDirector (Zhao et al., 2023) and our method do not need image conditions.
[…]
Qualitative results on few-shot.
Table 3. Quantitative results on few-shot. In human evaluation, the paired numbers show our method’s voting rate on the left, while previous methods’ voting rate on the right.
@@
Methods Text Align. $\uparrow$ Temporal Consist. $\uparrow$ Aesthetic Score $\uparrow$ ViCLIP Score $\uparrow$ VBench Average $\uparrow$ Appearance Diversity Motion Fidelity Temporal Consistency
LAMP (Wu et al., 2024) 26.19 87.89 18.41 22.68 67.41 59.1/40.9 63.6/36.4 60.6/39.4
DreamVideo (Wei et al., 2024) 27.48 90.52 19.17 25.10 71.75 66.7/33.3 53.0/47.0 50.0/50.0
MotionDirector (Zhao et al., 2023) 29.74 94.27 19.93 27.83 72.91 59.1/40.9 51.5/48.5 48.5/51.5
@@
(2) ${z}_{t-1}=\begin{cases}\frac{1}{\sqrt{\alpha_{t}}}\left({z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\epsilon_{\theta}({z}_{t},t)\right)+\sigma_{t},&t\leq\tau\\ \frac{1}{\sqrt{\alpha_{t}}}\left({z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\epsilon_{\theta,\theta_{K}}({z}_{t},t)\right)+\sigma_{t},&t>\tau\end{cases}$
where $\alpha$ is a constant controlling the noise strength.
4. Experiment
4.1. Experimental Setup
@@
Datasets. For evaluations on one video, we utilize LOVEU-TGVE-2023 dataset (Wu et al., 2023b). The dataset includes 76 reference videos, each accompanied by 4 prompts. Following  (Zhao et al., 2023), we design 3 additional prompts with significant appearance changes for each video to evaluate the diversity of the outputs. For evaluations on few-shot videos, we select 3-5 reference videos for each of the 13 categories in the UCF Sports dataset (Rodriguez et al., 2008; Soomro and Zamir, 2015). We design 8 prompts for the corresponding reference videos for each category. Similar to MotionDirector, we employ a large language model to create diverse prompts, each composed of [#Object] + [#Motion] + [#Appearance]. While the [#Object] remains consistent with the original prompts (for UCF data, [#Object] is the category name), the [#Motion] and [#Appearance] are very diverse.
@@
Baselines For evaluations on one-shot, we compare our method with the popular motion customization method, including MotionDirector (Zhao et al., 2023) and VMC (Jeong et al., 2024). In addition, we compare with controllable video generation methods, including VideoComposer (Wang et al., 2024), Control-A-Video (Chen et al., 2023), and Tune-a-Video (Wu et al., 2023a). Tune-a-Video, VMC, MotionDirector, and our method need fine-tuning, whereas VideoComposer and Control-A-Video do not need training. For evaluations on few-shot, we compare our approach with DreamVideo (Wei et al., 2024), LAMP (Wu et al., 2024), and MotionDirector.
@@
Metrics For automatic evaluations, we use three metrics: text alignment (Text Align.), temporal consistency (Temporal Consist.), and aesthetic score. Text alignment measures the average CLIP similarity (Radford et al., 2021) between each and the text prompt, reflecting how well the generated video aligns with the given description. Temporal consistency assesses the video’s coherence by CLIP similarity between frames. The aesthetic score is calculated using an aesthetic model (Kirstain et al., 2023). To provide a more comprehensive evaluation, we also leverage ViCLIP (Wang et al., 2023a) and VBench (Huang et al., 2024b). ViCLIP is used as a supplementary metric alongside TextCLIP. VBench is a benchmark that evaluates models across multiple dimensions. Due to space limitations, we report the average of these results (VBench Average), with detailed results provided in the appendix. For human evaluations, we invite 11 participants to complete a survey, resulting in a total of 1,419 responses. The survey assesses appearance diversity, motion fidelity, and temporal consistency. We show two videos: one generated by our method and the other by previous methods. For each evaluation dimension, users vote for the better video.
4.2. Evaluations on One-Shot
@@
Quantitative results. For automatic evaluations in Tab. 2, our method significantly improves text alignment and ViCLIP score, demonstrating that it suppresses appearance leakage. Our method also demonstrates superiority in other dimensions.
@@
Qualitative results. In the example on the left of Fig. 5, previous methods generate “grey sharks” exhibiting overfitting issue, while our method generates a realistic “sea dragon”. In the example on the right, our method generates a vivid “sandwich”, while the “sandwich” produced by other methods resembles the “pizza” from the reference video.
4.3. Evaluation on Few-Shot
@@
Quantitative results. As shown in Tab. 3, the automatic evaluation shows our method surpasses previous approaches across multiple dimensions. Human evaluations show that while our temporal consistency may not surpass that of DreamVideo and MotionDirector, our appearance diversity and motion fidelity are superior.
@@
Qualitative Results. Fig. 6 illustrates that MotionDirector occasionally exhibits appearance leakage from the training set. In the example on the left, DreamVideo generates a teddy bear holding an object resembling a suitcase and wearing a tie. It occurs because the training set includes humans carrying suitcases and wearing ties. MotionDirector generates a background featuring vertical reflective structures, similar to the background in the second reference video. In the right case, the backgrounds generated by previous methods lack a distinctly “cyberpunk” aesthetic and instead resemble those of reference videos. In contrast, our method avoids this issue.
4.4. Ablation
@@
Effect of Our Sub-Contributions. In Tab. 4, we evaluate the capabilities of each contribution: temporal attention purification (TAP), appearance highway (AH) and phased LoRA integration (PLI). TAP, AH, and PLI each significantly improve text alignment. AH improves text alignment and aesthetic score, but slightly compromises temporal consistency. However, combined with PLI, it not only further enhances text alignment and aesthetic score, but also mitigates the loss of motion consistency caused by AH.
Table 4. Effect of each component and effect of motion customization with different T2V-DM. We test on two popular T2V-DM models: ZeroScope(ZS) and ModelScope(MS).
Models
TAP
AH
PLI
Text Align. $\uparrow$
Temporal
Consist. $\uparrow$
Aesthetic
Score $\uparrow$
ViCLIP
Score $\uparrow$
VBench
Average $\uparrow$
@@
Effect of motion customization with different T2V-DM. Tab. 4 demonstrates that our method is effective across various foundation models using a spatial-temporal U-Net, such as ZeroScope (ZS) (Sterling, 2023) and ModelScope (MS) (Wang et al., 2023b).
Training with AH vs. Post-processing with AH. The AH can be used during both training and inference, or exclusively during inference. As shown in Tab. 5, the results are comparable. For simplicity, we use it with post-processing as our method.
Table 5. Training with AH vs. Post-processing with AH: (1) Training: apply it during both training and inference. (2) Post-processing: during training, the vanilla skip connection is used, while at inference, AH is employed. Experiments show that these two strategies are comparable.
Method
Text
Align. $\uparrow$
Temporal
Consist.
Aesthetic
Score $\uparrow$
ViCLIP
Score $\uparrow$
VBench
Average $\uparrow$
@@
Different choices of appearance highway’s scale. We scale the starting point of AH, i.e., hidden states of spatial transformer by a constant $\beta$ to strengthen its effect. As shown in Tab. 6, we change the $\beta$ and evaluate the model’s capabilities. When $\beta$ is small, increasing $\beta$ enhances the appearance information generated by T2V-DM and improves text alignment and aesthetic score. However, when $\beta$ becomes larger, further increases in $\beta$ lead to a decrease in these results. We hypothesize that this is due to the excessive magnitude of the original hidden states, which disrupts the data distribution and consequently degrades the generation quality. In addition, we found that enhancing AH leads to a decrease in temporal consistency. However, we mitigate this issue through temporal attention purification and phased LoRA integration.
Table 6. Different choices of appearance highway’s scale $\beta$.
Scale Text Align. $\uparrow$ Temporal Consist. $\uparrow$ Aesthetic Score $\uparrow$ ViCLIP Score $\uparrow$ VBench Average $\uparrow$
$\beta$=1 28.13 93.93 20.52 26.02 75.72
$\beta$=1.05 28.14 93.86 20.53 26.10 75.87
$\beta$=1.1 28.32 93.71 20.54 26.24 75.93
$\beta$=1.15 28.31 93.60 20.52 26.20 76.10
$\beta$=1.2 28.29 93.48 20.52 26.23 76.30
5. Conclusion
In this paper, we propose two novel strategies to separate motion from appearance during adapting T2V-DM to perform motion customization, including Temporal Attention Purification (TAP) and Appearance Highway (AH). In TAP, we utilize LoRA only to adapt the temporal attention, and in AH, we alter the starting point of skip connection from LoRA-adapted temporal transformer block to non-adapted spatial transformer block. Extensive experiments show that our method outperforms previous state-of-the-arts.
Acknowledgements.
This project is supported by National Natural Science Foundation of China under Grant 92370114.